OpenStudy (anonymous):

(4,0),(3,-1),(6,3),(2,-4) which are solutions to y=x-4 choose all correct answers (4,0) (3,-1) (6,3) (2,-4) medal to whoever helps and I will become a fan!!!

OpenStudy (anonymous):

@jhonyy9

OpenStudy (shinalcantara):

do you know x and y intercepts?

jhonyy9 (jhonyy9):

SO YOU KNOW THET y=x-4 just you need to try these sets like for (4,0) this mean thet x=4 need being y=0 0=4-4 this is correct --- true do you can continue it ?

OpenStudy (anonymous):

I honestly don't know a lot about algebra I just know if I don't turn in my homework I cant play in my game on Friday Im just trying to get it done..

jhonyy9 (jhonyy9):

so the second (3,-1) -1=3-4 -1=-1 true

jhonyy9 (jhonyy9):

(6,3) 3=6-4 3=2 false

OpenStudy (anonymous):

done?

jhonyy9 (jhonyy9):

the last one can you try it ?

jhonyy9 (jhonyy9):

(2,-4) -4 = 2-4

jhonyy9 (jhonyy9):

so how is it ?

OpenStudy (anonymous):

I think I got it but can you help me with two more questions?

jhonyy9 (jhonyy9):

you dont replied the last

OpenStudy (anonymous):

what?

jhonyy9 (jhonyy9):

-4=2-4

OpenStudy (anonymous):

false.

jhonyy9 (jhonyy9):

yes nice

OpenStudy (anonymous):

okay can you help me with the other questions?

jhonyy9 (jhonyy9):

write it on the question column please